### v1

@@ Abstract
The quality of Optical Music Recognition (OMR) systems is a rather difficult magnitude to measure. There is no lingua franca shared among OMR datasets that allows to compare systems’ performance on equal grounds, since most of them are specialised on certain approaches. As a result, most state-of-the-art works currently report metrics that cannot be compared directly. In this paper we identify the need of a common music representation language and propose the Music Tree Notation (MTN) format, thanks to which the definition of standard metrics is possible. This format represents music as a set of primitives that group together into higher-abstraction nodes, a compromise between the expression of fully graph-based and sequential notation formats. We have also developed a specific set of OMR metrics and a typeset score dataset as a proof of concept of this idea.
@@
Optical Music Recognition, Evaluation, Datasets, Computer Vision
@@ 1 Introduction
Written music has been part of the cultural heritage of humankind for many centuries. From neumes as old as a thousand of years to current-time western notation scores, people have built methods to make this most ephemeral of arts persistent through the passing of time. Rivers of ink have been imprinted on libraries worth of paper in order to preserve music along the ages, and much of it has indeed survived until today.
@@
It is therefore unsurprising that, given the sheer volume of notably interesting (and oftentimes, forgotten) pieces of music that have been endowed to our current generations, scholars have turned their attention to computers to aid them in the endeavour of preserving and analysing them. The task that concerns us in this work is Optical Music Recognition (OMR), which is the first and therefore fundamental step for the computational analysis of written scores; converting images or scans of music into a defined format a computer can process Calvo-Zaragoza et al (2021).
@@
Modern-day OMR has been greatly enriched with the advent of the powerful Deep Learning techniques developed during the last decade Calvo-Zaragoza et al (2021); Rebelo et al (2012). Nevertheless, the field itself remains fragmented, with few researchers fully devoted to it and each of them developing their own unique point of view and methodology Pacha (2018); Calvo-Zaragoza et al (2021). This is particularly evident when analysing the available datasets Pacha (2017), as can be seen in Table 1; most of them are restricted to specific steps or approaches Parada-Cabaleiro et al (2017); Tuggener et al (2018); Calvo-Zaragoza and Rizo (2018a) and almost none of them are compatible among each other (the most notable exception being the recent Shatri and Fazekas (2021), which incorporates its ground truth in multiple formats).
@@
One of the main points of disagreement among OMR researchers is the matter of the evaluation of models Hajič and Pecina (2017); Hajič jr. (2018); Mengarelli et al (2019); Calvo-Zaragoza et al (2021), which is but the visible consequence of many long-standing issues – some related to the inherent difficulty of the field (for an example, see Figure 1), but others related to the diversity of plausible approaches to music recognition.
@@
Figure 1: Fragment of a scan of a 1910 edition of Claude Debussy’s La Danse de Puck from the first volume of his Préludes. In this fragment, many of the reasons why modelling OMR scores is difficult can be seen – a two-staff system with voices that move along both of them, many small symbols concentrated in limited areas of the score, many independent voices sounding in unison, elements spanning multiple measures, chords, etc.
@@
Evaluation of OMR models is currently performed on a per-methodology basis Calvo-Zaragoza et al (2021); Rebelo et al (2012). If one considers the goal of OMR to produce a computational representation of music – whichever that final representation might be  [^1] Byrd et al. Byrd and Isaacson (2003) collected a very extensive and detailed set of requirements for a music notation system that is suitable for both score engraving and OMR, but to the best of our knowledge there is no recognition system that formally or explicitly applies them. – one would expect the metrics to obey a criterion of either some aspect of music understanding or quantifying the correctness of the final representation. Instead, most metrics revolve around measuring the fidelity of intermediate representations devoid of musical meaning, disregarding the full reconstruction of music scores altogether. Moreover, the benchmarks this evaluation metrics are computed upon are rarely widespread, with each methodology using their own. Commercial OMR systems have also been very vague about the specifics of their capacity.
@@
With the aim of setting a step forward towards unifying the efforts of the community, we propose a way to set a common framework for OMR. We believe the first step is to set a target final representation that accommodates a significant amount of use cases within Common Western Music Notation (CWMN). We explicitly avoid trying to cover the entirety of OMR, as the nature of notated music is diverse and heavily dependent on the time it was written and the instrument it was written for. Once this target representation is chosen, a way of easily working with it has to be developed. Currently used engraving formats are quite exhaustive, verbose and option-filled, which allows expressing the same score in a variety of ways – something that makes them quite impractical for direct comparison. Thus, we propose an intermediate representation that translates directly to the desired output but facilitates OMR operations – including analysis, evaluation and composition of results – regardless of the underlying approach. We propose doing so through a notation based on the idea of building a tree structure; leaves are the objects in the music score and intermediate nodes are their composition. The idea is exploiting the expression advantages of graph-based approaches while simplifying the kinds of relationships that such system can model. Incorporating a common notation format is also a way of unifying all existing datasets so that they can be both combined and compared on equal grounds, regardless of the methodology they were originally designed to work with.
@@
  • We try to bridge the gap between the different benchmark suites in OMR literature with a universal tree-based notation format designed to represent musical scores at the graphical level.
  • We also present a common evaluation toolkit which aims towards unify existing benchmark OMR tasks for fairer comparison.
  • We have produced a typeset dataset  [^2] The data shall be published upon acceptance alongside the conversion and evaluation code. using public domain works with permissive licenses.
@@ 2 Terminology
By structured output we refer to an output encoding format that allows the replication of the input score holistically and unambiguously. This is in contrast to partial outputs such as MIDI, which only encodes note pitch, duration and velocity and omits any other indication present on the score.
Another relevant point is the definition of an end-to-end OMR model. Modern day literature employs this term frequently, but it must be noted that it is used with a slightly different meaning than the term itself actually conveys. The most widely recognised definition of the OMR pipeline is perhaps that of the review penned by Rebelo et al. Rebelo et al (2012), which states four distinct steps:
@@
Most end-to-end approaches do not actually perform all four steps; rather, these models stop at step 3 with a notation that encodes most of the score’s information directly or indirectly, but cannot be directly engraved (sometimes requiring some additional interpretation before all music semantics can be successfully produced). Whenever we address end-to-end approaches, we refer to these. In this work we mostly address offline use cases of OMR – that is, recognition of raster images of scores. Nevertheless, we believe some of our ideas are equally applicable in online use cases Calvo-Zaragoza and Oncina (2014).
3 The Necessity for a Common OMR Evaluation Framework
In this section we provide thorough arguments as to why having a comprehensive and common OMR evaluation system is necessary, while also providing some insights into what we deem is the preferable approach given the current circumstances of the field.
@@ 3.1 OMR Needs a Standardised Output
Most OMR research today does not tackle the problem of converting a model’s output to a symbolic representation that can be engraved into a proper score. In their review of the field in 2021, Calvo-Zaragoza et al. Calvo-Zaragoza et al (2021) go as far as claiming that no research OMR system that they know of is actually capable of performing such a conversion [^3] There are however commercial OMR systemsBitteur (2004) that can perform this step., and since its publication we have only acknowledged one system that can Shishido et al (2021). One of their main justifications for this is the fact that most structured music engraving formats are either not mature enough to cover all possible use cases or simply impractical for OMR tasks. In practice, what this means is that very few OMR systems spawned in research contexts are employed by final users. Furthermore, it hinders OMR systems’ evaluation potential, as omitting music semantics inevitably reduces the level of interpretability of the models’ outputs.
Going back to the matter of whether notation systems are practical or mature enough for recognition, there are a few points that need to be made. If the final goal of recognition is a structured encoding that attempts to replicate the original score with the highest fidelity, there are currently two realistic choices: MusicXML Good (2001) and MEI Roland (2002). A point can be made about LilyPond’s The LilyPond Developement Team (2014) ability to engrave scores from text written in a fairly succinct and descriptive fashion, but since the specification of the format is subject to change and the fact that it under-specifies some aspects of the score for the sake of usability, we do not think it is realistic to use it as a comprehensive final output. Some effort has also been devoted into extending Humdrum’s **kern notation to make it capable of engraving scores uniquely Ríos-Vila et al (2023), but some of the same limitations as those seen for Lilypond apply.
MusicXML and MEI are somewhat similar in their base design ideas, with the former being more widely used within music score repository platforms and engraving software and the latter in libraries and academic-oriented settings. Perhaps the most obvious difference between the two is the fact that MusicXML is designed with the idea of enabling MIDI playback out-of-the-box at the cost of some redundancy, whereas MEI is more focused on musicological and archival applications. At the end of the day, they both can model the most widely used subset of CWMN music and they both have limitations derived from their use of XML and CWMN’s inherent flexibility; it is difficult to express co-depending symbols as a tree-like hierarchy. Therefore, both choices are probably equal in value in terms of expressiveness.
Thus, we believe it to be a net positive to make all OMR applications interoperable with these formats by design. The first advantage is making it possible to export scores back and forth and therefore increasing the amount of data available for recognition while making OMR more appealing for the end-user. The second one is that the set of requirements of an OMR system is grounded to those of the target application, making it simpler to cooperate among the community as everybody is talking the same language. Finally, and thanks to the fact that the targeted output is well defined, unambiguous metrics for end-to-end score recognition can be defined.
The big question is then how to make it practical for OMR practitioners to target these (rather complex) formats in a streamlined fashion. Our belief is that it is best not to use them directly, but rather incorporating a simplified intermediate format that makes it easy to convert to the structured outputs and also sets a semantic frame that adjusts well to the necessities of OMR.
3.2 On Computational Music Representations and their Tradeoffs
@@
The most flexible – and perhaps the most logical – computational representation of written music is most likely a graph, since it allows expressing $n$-dimensional arbitrary relationships between primitives and symbols. There exists a very widely known notation format based on this idea called Music Notation Graph (MuNG) introduced for the MUSCIMA++ Hajič and Pecina (2017). Some works have indeed succeeded in being able to reconstruct music structure from these graphs, such as Baró et al.’s Musigraph Baró et al (2022) or Garrido-Muñoz et al.’s Convolutional Recurrent Neural Networks Garrido-Munoz et al (2022). Nevertheless, to the best of our knowledge, there is no straightforward way of converting this graph output back into a structured encoding. Moreover, the cost of evaluation on graphs is high due to the complexity of matching graphs together.
At the other end of the spectrum lies the sequential music model, currently employed in some form or another by most end-to-end OMR approaches Calvo-Zaragoza et al (2017b); Calvo-Zaragoza and Rizo (2018b); Baró et al (2019); Alfaro-Contreras and Valero-Mas (2021); Ríos-Vila et al (2022b); Baró et al (2020); Torras et al (2021); Wen and Zhu (2022). These methods impose a relationship between the horizontal axis of the image input and the playback time of the score. However, this forces the insertion of music semantics in order to properly model specific musical concepts such as unison, as the only reliable way of aligning individual voices to their playback positions is by providing their starting times. This is, in fact, how ubiquitous formats such as MusicXML deal with this ambiguity – a “backup” token specifies that the internal clock of the engraver must be brought back a number of time partitions and start producing notes there. Current end-to-end OMR efforts are mostly restricted to either monophonic or homophonic music scores as a result.
This dichotomy of approaches defines a search space that is still being explored Alfaro-Contreras and Valero-Mas (2021); Alfaro-Contreras et al (2022); Baró et al (2019); Hajič and Pecina (2017). Our proposal is to try and exploit the advantages of both ends: the structural expresiveness of graphs and the ease of use of sequences. We propose modelling music as a tree structure, in which small semantic units (primitives) can be composed into complex high-level constructs (note groups, measures). The tree restriction makes it simpler to operate with, as the number of relationships is drastically reduced, relationship semantics are strictly hierarchical (and thus simpler to model) and its resemblance to the targeted formats makes conversion a simpler task. Also, imposing an ordering restriction within the tree as in a sequence makes matching computationally less demanding and allows modelling the idea of time and position somewhat implicitly.
3.3 The Complicated Endeavour of Music and Data
@@
Hajič and Pecina (2017)
MUSCIMA++
Handwritten Contemporary
Tuggener et al (2018, 2020)
DeepScores
Typeset
Tuggener et al (2023)
RealScores
Scanned
Shatri and Fazekas (2021)
DoReMi
Typeset
Parada-Cabaleiro et al (2017)
SEILS
Typeset Scanned Mensural
Baró et al (2020)
Baró Synthetic
Typeset
Baró et al (2020)
Pau Llinás
Handwritten Historical
Calvo-Zaragoza and Rizo (2018b)
PRIMuS
Typeset
Calvo-Zaragoza and Rizo (2018a)
Camera PRIMuS
Typeset
@@
One of the longest-standing and most crippling problems OMR has faced in the recent years is the lack of reliably annotated data for contexts other than recognition of typeset scores. Even in that case, the datasets are heavily opinionated and rarely incorporate structured music sources (see Table 1 for more detail).
In particular, there is a very noticeable lack of annotated data in CWMN set within historical contexts – that is, scores written from the 16th century up to the 20th. As can be seen in Table 1, only Baró et al.’s Baró et al (2020) Pau Llinás dataset is dedicated to music from this period. Ironically, some of the composers within it are among the most extensively studied musicians of all time: Bach, Beethoven, Mozart, Händel or Haydn, to name a few.
Given their importance to the art and craft of music, many efforts have been made to preserve their work. Some libraries have extensive repositories of historical music documents from a very broad period of time, some of which have devoted efforts specifically aimed at democratising access to big names in music noa (2023, 2008). Some of these original manuscripts are currently part of the public domain, which also enables them for use within OMR research.
The main question however is how to best incorporate this data into the field, as most of the original manuscripts do not have an associated ground truth that is useful for OMR. The most direct approach would be to collect a significant corpus of works and manually annotate it comprehensively (per-pixel symbol annotation, music score transcription, etc.). Nevertheless, this annotation is both extremely costly and time-consuming, and it entails an implicit choice of domain; scores that do not abide to the statistical properties of the collected corpus will still be intractable, and thus the manual annotation cycle would begin anew.
Therefore, our belief is that the way forward is not just about producing new hand-annotated datasets, but rather trying to exploit every single bit of data already publicly available. More specifically, we propose the creation of new OMR datasets based on existing transcriptions of public domain works. The goals are twofold: 1) to build a corpus of real scores soft-aligned (that is, a dataset where the position of measures within the page is known but the position of each individual object need not be) with existing transcriptions and 2) to enable the generation of old-looking synthetic scores, an avenue that is already being explored by Mayer et al. Mayer and Pecina (2021); Mayer (2022).
The downside of this approach is that producing symbol-level masks becomes considerably more complicated, which in practical terms shuts down the use of object detection-based methods in the short term. We nevertheless believe that this is probably going to be a temporary situation, as enabling recognition of these scores might potentially open the door for robust implicit segmentation or alignment systems with which an object-level dataset can be bootstrapped.
When it comes to the problem of alignment of scores, some issues do arise in the form of conflicting transcriptions and notation adaptations by editors or scholars. Nevertheless, heuristics can be set in place to detect consistent mistakes or usual deliberate changes – e.g. changes in the time signature, changes in the way voices are arranged on the staff, changes in clef and so on – and then apply the appropriate modifications back on the ground truth transcription. Those cases where the transcription is wrong are assumed to be minimal but possible nonetheless, as with any other human-produced ground truth.
Other roads for the use of non-annotated data currently exist to improve recognition efforts. In line with recent trends in the literature about building very large language models, many methods employ unlabelled input samples as some form of proxy task target. Examples of such techniques in Computer Vision related tasks include SimCLR Chen et al (2020) or Masked Autoencoders He et al (2022), which to the best of our knowledge have never been used in the context of OMR.

### v2

@@ Abstract
Modern-day Optical Music Recognition (OMR) is a fairly fragmented field. Most OMR approaches use datasets that are independent and incompatible between each other, making it difficult to both combine them and compare recognition systems built upon them. In this paper we identify the need of a common music representation language and propose the Music Tree Notation (MTN) format, with the idea to construct a common endpoint for OMR research that allows coordination, reuse of technology and fair evaluation of community efforts. This format represents music as a set of primitives that group together into higher-abstraction nodes, a compromise between the expression of fully graph-based and sequential notation formats. We have also developed a specific set of OMR metrics and a typeset score dataset as a proof of concept of this idea.
@@
Optical Music Recognition, Representation, Evaluation, Datasets, Computer Vision
@@ 1 Introduction
Written music has been part of the cultural heritage of humankind for many centuries. From neumes as old as a thousand years to current-time Western notation scores, people have built methods to make this most ephemeral of arts persistent over time. Rivers of ink have been imprinted on libraries worth of paper to preserve music along the ages, and much of it has indeed survived until today.
@@
It is therefore unsurprising that, given the sheer volume of notably interesting (and oftentimes, forgotten) pieces of music that have been endowed to our current generations, scholars have turned their attention to computers to aid them in the endeavour of preserving and analysing them. The task that concerns us in this work is Optical Music Recognition (OMR), which is fundamental step for the computational analysis of written scores; converting images or scans of music into a defined format a computer can process Calvo-Zaragoza et al (2021).
@@
Modern-day OMR has been greatly enriched with the advent of the powerful deep-learning techniques developed during the last decade Calvo-Zaragoza et al (2021); Rebelo et al (2012). Nevertheless, the field itself remains fragmented, with few researchers fully devoted to it and each of them developing their unique point of view and methodology Pacha (2018); Calvo-Zaragoza et al (2021). This is particularly evident when analysing the available datasets Pacha (2017), as can be seen in Table 1; most of them are restricted to specific steps or approaches Parada-Cabaleiro et al (2017); Tuggener et al (2018); Calvo-Zaragoza and Rizo (2018a) and almost none of them are compatible with each other (the most notable exception being the DoReMi dataset introduced in 2021 Shatri and Fazekas (2021), which incorporates its ground truth in multiple formats).
@@
Another point of disagreement among OMR researchers is the matter of the evaluation of models Hajič and Pecina (2017); Hajič jr. (2018); Mengarelli et al (2019); Calvo-Zaragoza et al (2021), which is but the visible consequence of many long-standing issues – some related to the inherent difficulty of the field (for an example, see Figure 1), but others related to the diversity of plausible approaches to music recognition.
@@
Figure 1: Fragment of a scan of a 1910 edition of Claude Debussy’s La Danse de Puck from the first volume of his Préludes. In this fragment, many of the reasons why modelling OMR scores is difficult can be seen – a two-staff system with voices that move along both of them, certain areas of the score accumulating many small symbols, many independent voices sounding in unison, elements spanning multiple measures, chords, etc.
@@
Evaluation of OMR models is currently performed on a per-methodology basis Calvo-Zaragoza et al (2021); Rebelo et al (2012). If one considers the goal of OMR to produce a computational representation of music – whichever that final representation might be  [^1] Byrd et al. Byrd and Isaacson (2003) collected a very extensive and detailed set of requirements for a music notation system that is suitable for both score engraving and OMR, but to the best of our knowledge there is no recognition system that formally or explicitly applies them. – one would expect the metrics to obey a criterion of either some aspect of music understanding or quantifying the correctness of the final representation. Instead, most metrics revolve around measuring the fidelity of intermediate representations devoid of musical meaning (e.g. sequences of tokens or bounding boxes), disregarding the full reconstruction of music scores altogether. Moreover, the benchmarks these evaluation metrics are computed upon are rarely widespread, with each methodology using their own. Commercial OMR systems have also been very vague about the specifics of their capacity.
With the aim of setting a step forward towards unifying the efforts of the community, we propose a way to set a shared framework for OMR. We believe the first step is to set a target final representation that accommodates a significant amount of use cases within Common Western Music Notation (CWMN). This is the music notation system that has been employed in Europe from the early 18th century until today.
Focusing on CWMN specifically instead of the broad range of western notations related to it is not an arbitrary choice. While CWMN did indeed evolve from Mensural notation, which was employed from the 13th to the 17th century, and consequentially shares a considerable amount of graphical constructs, this system has semantic concepts that make them quite distinct. For instance, the concept of beat as we know it today was developed during the first half of the 18th century; before that, the rules that described the relation of note types with their divisions implied different abstractions and note types that are not compatible with those of modern notation. Moreover, there are constructs such as ligatures that have very unique properties but have disappeared from CWMN. In this regard, given that each notation system has a distinct nature and given the necessity for OMR systems in CWMN, we believe it is best to commit to this specific notation alone.
@@
Once this target representation is chosen, a way of easily working with it has to be developed. Currently used encoding formats are quite exhaustive, verbose and option-filled, which allows expressing the same score in a variety of ways – something that makes them impractical for both recognition and direct comparison. Thus, we propose an intermediate representation that translates directly to the desired output and offers analysis, conversion and evaluation tools regardless of the underlying approach. We propose doing so through a notation based on the idea of building a tree structure; leaves are the objects in the music score and intermediate nodes are their combination. The idea is exploiting the structural advantages of graph-based approaches over sequence-based ones while restricting the algorithmic complexity of working with arbitrary graphs. Incorporating a shared notation format is also an opportunity for unifying all existing datasets so that they can be both combined and compared on equal grounds, regardless of the methodology they were originally designed to work with.
@@
  • We try to bridge the gap between the different benchmark suites in OMR literature with a universal tree-based notation format designed to represent musical scores at the graphical level [^2] Repository of the project: https://github.com/CVC-DAG/comref-converter.
  • We also present an evaluation toolkit which aims towards unify existing benchmark OMR tasks for fairer comparison.
  • We have produced a typeset dataset using public domain works with permissive licenses  [^3] Link to the dataset https://datasets.cvc.uab.cat/comref/comref.zip.
@@ 2 Terminology
By structured output we refer to an output encoding format that allows the replication of the input score exhaustively and unambiguously. This is in contrast to formats such as MIDI, that only encode certain aspects of music (in this case, playback information).
Another relevant point is the definition of an end-to-end OMR model. Modern day literature employs this term frequently, but it must be noted that it is used with a slightly different meaning than the term itself actually conveys. The most widely recognised definition of the OMR pipeline is perhaps that of the review authored by Rebelo et al. Rebelo et al (2012), which states four distinct steps:
@@
Most end-to-end approaches do not actually perform all four steps; rather, these models stop at step 3 with a notation that encodes most of the score’s information directly or indirectly, but cannot be directly engraved (sometimes requiring an additional processing step before all music semantics can be successfully generated). Whenever we address end-to-end approaches, we refer to these. In this work we mostly address offline use cases of OMR – that is, recognition of raster images of scores. Nevertheless, we believe some of our ideas are equally applicable to online use cases Calvo-Zaragoza and Oncina (2014).
3 The Necessity for an OMR Framework
In this section we comment on some of the main issues of the field of OMR and provide arguments as to why having a unified OMR framework is useful and necessary.
@@ 3.1 OMR Needs a Standardised Output
Most OMR research today does not tackle the problem of converting a model’s output to a symbolic representation that can be engraved into a proper score. In their review of the field in 2021, Calvo-Zaragoza et al. Calvo-Zaragoza et al (2021) go as far as claiming that no research OMR system that they know of is actually capable of performing such a conversion [^4] There are however commercial OMR systemsBitteur (2004) that can perform this step., and since its publication we have only acknowledged one system that can Shishido et al (2021). One of their main justifications for this is the fact that most structured music engraving formats are either not mature enough to cover all possible use cases or impractical for OMR tasks due to their complexity.
There are currently two main choices for music score exchange formats: MusicXML Good (2001), devised by Recordare and now managed by a W3C committee and the Music Encoding Initiative (MEI), designed by Perry Roland Roland (2002). Other formats such as Lilypond The LilyPond Developement Team (2014) (an open-source text-based music engraving software) or Humdrum **kern Huron (20??) (a notation format for the automatic musical analysis tools in the Humdrum suite) have been used for OMR as they allow engraving scores very succinctly. Nevertheless, these formats are not designed to represent scores precisely and unambiguously – beaming groups are not explicitly modelled, multiple representations of the same score through permutations are allowed (although **kern has been extended to fix this Ríos-Vila et al (2023)), ornaments, expression signs, dynamics and other non-structural elements are only partially supported, etc. As a consequence, these formats are not suited to engrave exact replicas of any music score.
MusicXML and MEI are somewhat similar in their base design ideas, with the former being more widely used within music score repository platforms and engraving software and the latter in archival settings. Thus, we believe it is important to make all OMR applications interoperable with these formats by design. Doing so makes it easier to collect data for OMR applications and also simplifies deploying OMR systems. Having a well-defined target representation has the additional side-effect of fixing the requirements for what is expected from an OMR system, encouraging joining efforts within the community.
Since working with these formats directly can be difficult, it would be interesting to have a different intermediate representation that acknowledges the requirements of OMR but can convert to other formats when required.
3.2 Trade-offs for Structured Representations of Music for OMR
To this day, there is no straightforward answer as to what is the best basic data structure to represent music for OMR purposes. There are various approaches being explored in the community, each having their own set of trade-offs.
@@ 3.2 Trade-offs for Structured Representations of Music for OMR
The most flexible – and perhaps the most logical – computational representation of written music is most likely a graph, since it allows expressing arbitrary relationships between primitives and symbols. There exists a very widely known notation format based on this idea called Music Notation Graph (MuNG) introduced for the MUSCIMA++ Dataset Hajič and Pecina (2017). Some works have indeed succeeded in being able to reconstruct music structure from these graphs, such as Baró et al.’s Musigraph Baró et al (2022) or Garrido-Muñoz et al.’s Convolutional Recurrent Neural Networks Garrido-Munoz et al (2022). Nevertheless, to the best of our knowledge, there is no straightforward way of converting this graph output back into a structured encoding. Moreover, the cost of evaluation on graphs is high due to the complexity of matching them.
At the other end of the spectrum lies the sequential music model, currently employed in some form or another by most end-to-end OMR approaches Calvo-Zaragoza et al (2017b); Calvo-Zaragoza and Rizo (2018b); Baró et al (2019); Alfaro-Contreras and Valero-Mas (2021); Alfaro-Contreras et al (2022); Ríos-Vila et al (2022b); Baró et al (2019, 2020); Torras et al (2021); Wen and Zhu (2022); Edirisooriya et al (2021); Ríos-Vila et al (2023). These methods impose a relationship between the horizontal axis of the image input and the playback time of the score. However, this forces either having to model music time semantics when inferring the structure of the score or introducing simplifications to music notation. The latter is the most common, with information such as stem directions or beaming groups being abstracted away to varying degrees.
3.3 Annotated Data is Scarce
@@
Hajič and Pecina (2017) MUSCIMA++ Handwritten Contemporary
Tuggener et al (2018, 2020) DeepScores Typeset
Tuggener et al (2023) RealScores Scanned
Shatri and Fazekas (2021) DoReMi Typeset
Parada-Cabaleiro et al (2017) SEILS Typeset Scanned Mensural
Baró et al (2020) Baró Synthetic Typeset
Baró et al (2020) Pau Llinás Handwritten Historical
Calvo-Zaragoza and Rizo (2018b) PRIMuS Typeset
Calvo-Zaragoza and Rizo (2018a) Camera PRIMuS Typeset
@@
One of the longest-standing problems in OMR is the lack of reliably annotated data for contexts other than recognition of typeset scores. Even in that case, the datasets are usually tailored to specific recognition techniques and rarely incorporate structured music sources (see Table 1 for more detail). Most datasets are incompatible between each other as a result.
In this sense, it would be very positive that all datasets provide an additional structured source in a common format in a way similar to Shatri and Fazekas (2021), so that the original data can be used in contexts outside the ones initially intended. This would help develop better systems overall by having the possibility of ensembling datasets in those domains where there are many of them available.
However, there are some domains for which there is barely any data available. An example of this are historical CWMN, written from between the 17th and 20th centuries. As can be seen in Table 1, only Baró et al.’s Baró et al (2020) Pau Llinás dataset is dedicated to music from this period. This is caused for various reasons:
  • The only realistic way of annotating a corpus of handwritten music from scratch given the tools existing today is doing so manually, which is extremely time-consuming.
  • Crowdsourcing annotation of music scores is difficult because it requires trained individuals. Some music scores in this period are notoriously difficult to read.
Making it possible to incorporate structured sources opens up the possibility of using existing public domain transcriptions works of widely known authors and trying to synchronise them to the original public manuscripts (e.g. Bach noa (2008) or Beethoven noa (2023)) at a significantly lower cost.